# Freeness of arrangements of lines and one conic with ordinary quasi-homogeneous singularities

### v4

@@ Abstract
The main purpose of the present paper is to provide a partial classification, performed with respect the weak-combinatorics, of arrangements consisting of lines and one smooth conic with quasi-homogeneous ordinary singularities that are free.
@@ 1 Introduction
The present paper is devoted to arrangements of lines and exactly one conic in the complex projective plane with quasi-homogeneous ordinary singularities. Our main motivation comes from a very active area of research devoted to free arrangements of rational curve arrangements in the plane and the so called Numerical Terao’s Conjecture which focuses on the so-called weak combinatorics of a given arrangement.
@@
Let $C=\{C_{1},...,C_{k}\}\subset\mathbb{P}^{2}_{\mathbb{C}}$ be a reduced curve such that each irreducible component $C_{i}$ is smooth. The weak combinatorics of $C$ is a vector of the form $(d_{1},...,d_{s};t_{1},...,t_{p})$, where $d_{i}$ denotes the number of irreducible components of $C$ of degree $i$, and $t_{j}$ denotes the number of singular points of a curve $C$ of a given analytic type $T_{j}$.
@@
This conjecture is somehow a natural generalization of the classical Terao’s conjecture on (central) hyperplane arrangements, where we focus on the intersection posets of arrangements as the decisive objects. It is worth recalling that the author with Alexandru Dimca showed that if $\mathcal{CL}$ is an arrangement of $k\geqslant 1$ smooth conics and $d\geqslant 1$ lines that admits nodes, tacnodes, and ordinary triple points as singularities, that the Numerical Terao’s Conjecture holds for this class of curves [5]. On the other hand, Marchesi and Vallés in [10] gave a counterexample to the Numerical Terao’s Conjecture in the class of line arrangements or, more precisely, in the class of triangular line arrangements.
The main aim of the present paper is to study the freeness of arrangements consisting of $d\geqslant 3$ lines and exactly one smooth conic that admit quasi-homogeneous ordinary singularities, but our motivation comes from a completely different perspective that we want to explain now.
We have several interesting invariants that can be attached to the Jacobian ideal associated with a defining equation $f\in S:=\mathbb{C}[x,y,z]$ of a reduced plane curve $C\,:f=0$. One such invariant is the minimal degree of Jacobian relations, which is defined as the minimal degree of a non-trivial triple $(a,b,c)\in S^{3}$ satisfying the condition that
@@
In this context, it is worth recalling a pair of non-free arrangements of $d=9$ lines constructed by Ziegler in [13] with the property that the arrangements have the same intersection lattices, so they have the same weak combinatorics which is $(d,t_{2},t_{3})=(9;18,6)$, but they have different minimal degrees of the Jacobian relations. These two line arrangements are distinguished by the property that in one case the $6$ triple points are on a smooth conic, and in the other case they are not. From our point of view, it is natural to include this ghostly existing conic passing through $6$ points and then to study the homological properties of the resulting conic-line arrangement. In the present paper, motivated by Ziegler’s example and a recent paper by Dimca and Sticlaru [7], we want to study the freeness of arrangements consisting of one smooth conic and $d\geqslant 3$ lines admitting ordinary quasi-homogeneous singularites, in the hope of better understanding the freeness property from the perspective of weak combinatorics and the minimal degree of Jacobian relations. This setting allows to provide a detailed partial classification result on admissible weak combinatorics of free arrangements with $3\leqslant d\leqslant 10$ lines and one conic having $n_{2}$ nodes, $n_{3}$ ordinary triple, and $n_{4}$ ordinary quadruple points as singularities (so singularities of types $A_{1},D_{4}$, and $X_{9}$ according to Arnold’s classification [1]). Recall that an ordinary singularity of multiplicity $m$ is quasi-homogeneous if $m<5$, see [4, Exercise 7.31].
@@
In order to decide whether a certain weak combinatorics can be realized over the real or complex numbers, one can determine (presumably effective) numerical constraints, such as Hirzebruch-type inequalities. Here we present a general tool that we can apply to reduced plane curves with ordinary double, triple, and quadruple points.
@@
Let $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ be a reduced plane curve of degree $d\geqslant 6$ admitting only $n_{2}$ nodes, $n_{3}$ ordinary triple and $n_{4}$ ordinary quadruple points. Then one has
$9n_{2}+\frac{117}{4}n_{3}+60n_{4}\leqslant 5d^{2}-6d.$
In the paper we work over the complex numbers and our symbolic computations are preformed using SINGULAR [3].
@@ 2 Preliminaries
$\mu_{p}=\dim_{\mathbb{C}}\left(\mathbb{C}[x,y]/\bigg{\langle}\frac{\partial f}{\partial x},\frac{\partial f}{\partial y}\bigg{\rangle}\right)$
@@
$\tau_{p}=\dim_{\mathbb{C}}\left(\mathbb{C}[x,y]/\bigg{\langle}f,\frac{\partial f}{\partial x},\frac{\partial f}{\partial y}\bigg{\rangle}\right)$
@@
Moreover, if $C\,:f=0$ is a reduced plane curve with only quasi-homogeneous singularities, then
@@
which means that the total Tjurina number of $C$ is equal to the total Milnor number of $C$.
@@
Now we are in a position to define the freeness in the language of the minimal degree of (non-trivial) Jacobian relations and the total Tjurina numbers following [8].
@@
Let $C\,:f=0$ be a reduced curve in $\mathbb{P}^{2}_{\mathbb{C}}$. Then the curve $C$ with $r:={\rm mdr}(f)\leqslant(d-1)/2$ is free if and only if
$(d-1)^{2}-r(d-r-1)=\tau(C).$ (1)
In order to perform our classification, we will need the following result [6, Theorem 2.1].
@@
Let $C\,:\,f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}_{\mathbb{C}}$ having only quasi-homogeneous singularities. Then
@@
It is worth recalling that the Arnold exponent of a given reduced curve $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ is defined as the minimum over all Arnold exponents of singular points $p$ in $C$. In modern language, the Arnold exponents of singular points are nothing else but the log canonical thresholds of singularities ${\rm lct}_{p}$. In the case of ordinary singularities, we have the following result [2, Theorem 1.3].
@@
If $p=(0,0)\in\mathbb{C}^{2}$ is an ordinary singularity of multiplicity $r$ determined by $C\,:\,f=0$, then ${\rm lct}_{p}(f)=\frac{2}{r}$.
@@ 3 Partial Classification
Our classification procedure is based on the following general approach that is based on combinatorial constraints that comes from the combinatorics of free curves.
Recall that if $\mathcal{CL}\subset\mathbb{P}^{2}_{\mathbb{C}}$ is an arrangement of $d\geqslant 3$ lines and one conic having $n_{2}$ nodes, $n_{3}$ ordinary triple and $n_{4}$ ordinary quadruple points, then
@@
and this is what we call as a naive count.
The second constraint concerns the total Tjurina number of $\mathcal{CL}\,:f=0$ which has degree $d+2$ and has $r:={\rm mdr}(f)$, namely
@@
so our problem here boils down to finding constraints on $r={\rm mdr}(f)$. We have the following result.
@@
Let $\mathcal{CL}\,f=0$ be an arrangement of $d\geqslant 3$ lines and one smooth conic that admits only ordinary singularities of multiplicity $<5$. Assume that $\mathcal{CL}$ is free, then
@@
Obviously ${\rm mdr}(f)\leqslant\lfloor\frac{d+1}{2}\rfloor$. Moreover, if $\mathcal{CL}$ admits only ordinary singularities with multiplicities $<5$, then $\alpha_{\mathcal{CL}}=\frac{1}{2}$, and this follows from Remark 2.6. This gives us that
@@
For $d\geqslant 3$ and $r:=r(d)\in\bigg{\{}\bigg{\lceil}\frac{d-2}{2}\bigg{\rceil},\bigg{\lfloor}\frac{d+1}{2}\bigg{\rfloor}\bigg{\}}$, we consider the following Diophantine system of equations:
@@
Then for admissible values of $d$ and $r(d)$, we find all non-negative integer solutions $(n_{2},n_{3},n_{4})$ obtaining a complete weak combinatorial description of the expected free arrangements. The last step boils down to deciding on the existence/non-existence of a geometric realization of a given weak combinatorics, which is a completely non-trivial problem. In our classification we use a well-known (even folkloric) result which tells us that all loopless matroids of rank $3$ with up to $6$ elements are vector matroids which can be represented geometrically as line arrangements in $\mathbb{P}^{2}_{\mathbb{K}}$ with $\mathbb{K}$ being any infinite field.
@@
    Since $\tau(\mathcal{CL}_{1})=12$, because $(n_{2},n_{3},n_{4})=(3,0,1)$, and ${\rm mdr}(Q_{1})=2$, hence $\mathcal{CL}_{1}$ is free.
@@
    Since $(n_{2},n_{3},n_{4})=(0,3,0)$, $\tau(\mathcal{CL}_{2})=12$ and ${\rm mdr}(Q_{2})=2$, hence $\mathcal{CL}_{2}$ is free.
@@
    Observe that $\tau(\mathcal{CL}_{3})=19$, since we have $(n_{2},n_{3},n_{4})=(2,2,1)$, and ${\rm mdr}(Q_{3})=2$, hence $\mathcal{CL}_{3}$ is free.
@@
    Since $\tau(\mathcal{CL}_{4})=28$, because we get the required intersections, and ${\rm mdr}(Q_{4})=2$, hence $\mathcal{CL}_{4}$ is free.
@@
    Observe that $\tau(\mathcal{CL}_{5})=27$ and ${\rm mdr}(Q_{5})=3$, hence $\mathcal{CL}_{5}$ is free.
@@
    Observe that $\tau(\mathcal{CL}_{6})=39$, since $(n_{2},n_{3},n_{4})=(3,0,4)$, and ${\rm mdr}(Q_{6})=2$, hence $\mathcal{CL}_{6}$ is free.
@@
    Observe that $\tau(\mathcal{CL}_{7})=37$, since $(n_{2},n_{3},n_{4})=(3,4,2)$, and ${\rm mdr}(Q_{7})=3$, hence $\mathcal{CL}_{7}$ is free.
@@
    Observe that $\tau(\mathcal{CL}_{8})=37$, since $(n_{2},n_{3},n_{4})=(6,1,3)$, and ${\rm mdr}(Q_{8})=3$, hence $\mathcal{CL}_{8}$ is free.
    To complete our classification for $d=6$, we need to show that the weak combinatorics $(n_{2},n_{3},n_{4})\in\{(0,3,3),(0,7,1)\}$ cannot be realized geometrically over the reals. By reasoning as in the previous situation, in both cases we get at a contradiction by arriving at a situation where two lines in a given arrangement intersect at two (or more) different points.
@@
    We can easily observe that $\tau(\mathcal{CL}_{9})=49$ and ${\rm mdr}(Q_{9})=3$, hence $\mathcal{CL}_{9}$ is free.
@@
    One can easily check that $\tau(\mathcal{CL}_{10})=48$ and ${\rm mdr}(Q_{10})=4$, hence $\mathcal{CL}_{10}$ is free.
@@
    One can check that $\tau(\mathcal{CL}_{11})=48$ and ${\rm mdr}(Q_{10})=4$, hence $\mathcal{CL}_{11}$ is free.
@@
    We can check that $(n_{2},n_{3},n_{4})=(5,5,4)$, so we have $\tau(\mathcal{CL}_{12})=61$, and ${\rm mdr}(Q_{12})=4$, hence $\mathcal{CL}_{12}$ is free.
@@
    We can check that $(n_{2},n_{3},n_{4})=(2,8,3)$, so we have $\tau(\mathcal{CL}_{13})=61$, and ${\rm mdr}(Q_{13})=4$, hence $\mathcal{CL}_{13}$ is free.
@@
    We can check that $(n_{2},n_{3},n_{4})=(8,2,5)$, so we have $\tau(\mathcal{CL}_{14})=61$, and ${\rm mdr}(Q_{14})=4$, hence $\mathcal{CL}_{14}$ is free.
@@
    Since $\tau(\mathcal{CL}_{15})=76$ and ${\rm mdr}(\mathcal{CL}_{15})=4$, hence $\mathcal{CL}_{15}$ is free.
@@
    Since $\tau(\mathcal{CL}_{16})=93$ and ${\rm mdr}(Q_{16})=4$, hence $\mathcal{CL}_{16}$ is free.

### v5

@@ Abstract
The main purpose of the present paper is to provide a partial classification, performed with respect the weak-combinatorics, of free arrangements consisting of lines and one smooth conic with quasi-homogeneous ordinary singularities.
@@ 1 Introduction
The present paper is devoted to arrangements of lines and exactly one conic in the complex projective plane with quasi-homogeneous ordinary singularities. Here by an arrangement we understand a reduced and reducible plane curve while by a reduced curve we mean a curve that might be also irreducible. Our main motivation comes from a very active area of research devoted to free arrangements of rational curve arrangements, where an arrangement is free if its associated module of derivations is a free module over the coordinate ring of the plane, and the so-called Numerical Terao’s Conjecture which focuses on the so-called weak combinatorics of a given arrangement.
@@
Let $C=\{C_{1},...,C_{k}\}\subset\mathbb{P}^{2}_{\mathbb{C}}$ be a reduced curve such that each irreducible component $C_{i}$ is smooth. The weak combinatorics of $C$ is a vector of the form $(d_{1},...,d_{s};m_{1},...,m_{p})$, where $d_{i}$ denotes the number of irreducible components of $C$ of degree $i$, and $m_{j}$ denotes the number of singular points of a curve $C$ of a given type $M_{j}$.
In the above definition we refer to types of singularities, which for us are determined by their local normal forms. In our paper we will use Arnold’s classification of local normal forms presented in [1].
@@
This conjecture is somehow a natural generalization of the classical Terao’s conjecture on (central) hyperplane arrangements, where we focus on the intersection posets of arrangements as the decisive objects. It is worth recalling that the author with Alexandru Dimca showed that if $\mathcal{CL}$ is an arrangement of $k\geqslant 1$ smooth conics and $d\geqslant 1$ lines that admits nodes, tacnodes, and ordinary triple points as singularities, that the Numerical Terao’s Conjecture holds for this class of curves [6]. On the other hand, Marchesi and Vallés in [13] gave a counterexample to the Numerical Terao’s Conjecture in the class of line arrangements or, more precisely, in the class of triangular line arrangements. Very recently, the author with Luber and Kühne in [12] have shown that Numerical Terao’s Conjecture holds for arrangements with up to $d=12$ lines, and there are some counterexamples to this conjecture for $d=13$ lines. Based on that result, it is somehow natural to ask whether we can find a counterexample to Numerical Terao’s Conjecture in the least general setting comparing with the line arrangement scenario. More precisely, we seek for the simplest possible counterexample, if existing, in the setting of smooth rational curve arrangements with ordinary quasi-homogeneous singularities, so we want to work with conic-line arrangements admitting ordinary points of multiplicity $2,3$ and $4$. For the completeness of our presentation we recall that an ordinary singularity of multiplicity $m$ is quasi-homogeneous if $m<5$, and this can be found in [5, Exercise 7.31]. Note also that our assumption of working with quasi-homogeneous singularities follows from the fact that, in general, singularities of conic-line arrangements are not quasi-homogeneous, and this leads to some pathologies as indicated in [3, Example 4.6]. Another motivation for our studies comes from a completely different perspective that we want to explain now. We have several interesting invariants that can be attached to the Jacobian ideal associated with a defining equation $f\in S:=\mathbb{C}[x,y,z]$ of a reduced plane curve $C\,:f=0$. One such invariant is the minimal degree of Jacobian relations, which is defined as the minimal degree of a non-trivial triple $(a,b,c)\in S^{3}$ satisfying the condition that
@@
In this context, it is worth recalling a pair of non-free arrangements of $d=9$ lines constructed by Ziegler in [17] with the property that the arrangements have the same intersection lattices, so they have the same weak combinatorics which is $(d,t_{2},t_{3})=(9;18,6)$, but they have different minimal degrees of the Jacobian relations. These two line arrangements are distinguished by the property that in one case the $6$ triple points are on a smooth conic, and in the other case they are not. From our point of view, it is natural to include this ghostly existing conic passing through $6$ points and then to study the homological properties of the resulting conic-line arrangement. In the present paper, motivated by Ziegler’s example and a recent paper by Dimca and Sticlaru [8], we want to study the freeness of arrangements consisting of one smooth conic and $d\geqslant 3$ lines admitting ordinary quasi-homogeneous singularites, in the hope of better understanding the freeness property from the perspective of weak combinatorics and the minimal degree of Jacobian relations. This setting allows to provide a detailed partial classification result on admissible weak combinatorics of free arrangements with $3\leqslant d\leqslant 10$ lines and one conic having $n_{2}$ nodes, $n_{3}$ ordinary triple, and $n_{4}$ ordinary quadruple points as singularities (i.e. singularities of types $A_{1},D_{4}$, and $X_{9}$ according to Arnold’s classification [1]). Since we are working with only one conic, we abbreviate the representation of the vectors of weak combinatorics to the form $(d;n_{2},n_{3},n_{4})$.
@@
In order to decide whether a certain weak combinatorics can be realized over the real or complex numbers, one can determine (presumably effective) numerical constraints, such as Hirzebruch-type inequalities. Here we present a general tool that we can apply to reduced plane curves with ordinary double, ordinary triple and ordinary quadruple points.
@@
Let $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ be a reduced plane curve of degree $m\geqslant 6$ admitting only $n_{2}$ ordinary double, $n_{3}$ ordinary triple and $n_{4}$ ordinary quadruple points. Then one has
$9n_{2}+\frac{117}{4}n_{3}+60n_{4}\leqslant 5m^{2}-6m.$
In the paper we work, most of the time, over the complex numbers and our symbolic computations are preformed using SINGULAR [4].
@@ 2 Preliminaries
$\mu_{p}=\dim_{\mathbb{C}}\left(\mathbb{C}\{x,y\}/\bigg{\langle}\partial_{x}\,f,\partial_{y}\,f\bigg{\rangle}\right)$
@@
$\tau_{p}=\dim_{\mathbb{C}}\left(\mathbb{C}\{x,y\}/\bigg{\langle}f,\partial_{x}\,f,\partial_{y}\,f\bigg{\rangle}\right)$
@@
Recall that a singularity is called quasi-homogeneous if and only if there exists a holomorphic change of variables so that the defining equation becomes weighted homogeneous. If $C=\{f=0\}$ is a reduced plane curve with only quasi-homogeneous singularities, then by [16, Satz] one has
@@
which means that the total Tjurina number is equal to the total Milnor number of $C$.
@@
A reduced curve $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ is free if the Jacobian ideal $J_{f}$ is saturated with respect to $\mathfrak{m}=\langle x,y,z\rangle$.
It is somewhat difficult to check the freeness property using the above definition. However, it turns out that we can check the freeness property using the language of the minimal degree of (non-trivial) Jacobian relations and the total Tjurina numbers according to a result due to du Plessis and Wall [9].
Theorem 2.4.
Let $C=\{f=0\}$ be a reduced curve in $\mathbb{P}^{2}_{\mathbb{C}}$. One has
$(d-1)^{2}-r(d-r-1)=\tau(C)$ (1)
if and only if $C=\{f=0\}$ is a free curve, and then $r\leqslant(d-1)/2$.
In order to perform our classification, we will need the following result [7, Theorem 2.1].
@@
Let $C=\{f=0\}$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}_{\mathbb{C}}$ having only quasi-homogeneous singularities. Then
@@
It is worth recalling that the Arnold exponent of a given reduced curve $C\subset\mathbb{P}^{2}_{\mathbb{C}}$ is defined as the minimum over all Arnold exponents of singular points $p$ in $C$. In modern language, the Arnold exponents of a singular point $p$ of a curve $C$ are nothing but the log canonical thresholds ${\rm lct}_{p}$ of the pair $(C,p)$. In the case of ordinary singularities, we have the following result [2, Theorem 1.3].
@@
If $p=(0,0)\in\mathbb{C}^{2}$ is an ordinary singularity of multiplicity $r$ determined by $C=\{f=0\}$, then ${\rm lct}_{p}(f)=\frac{2}{r}$.
@@ 3 Partial Classification
Our classification procedure is based on the following general approach which uses combinatorial constraints derived from the combinatorics of free curves.
If $\mathcal{CL}\subset\mathbb{P}^{2}_{\mathbb{C}}$ is an arrangement of $d\geqslant 3$ lines and one conic having $n_{2}$ nodes, $n_{3}$ ordinary triple and $n_{4}$ ordinary quadruple points, then by Bézout’s theorem we have
@@
and this is what we call the naive count. This simple combinatorial count is crucial because it allows us to determine which weak combinatorics are admissible for our conic-line arrangements. For instance, there does not exist an arrangement of two lines and one conics with $6$ ordinary double intersections. However, this combinatorial count is not strong enough for our purposes and we need to find additional restrictions. In this sense, we introduce the second constraint which concerns the total Tjurina number of $\mathcal{CL}=\{f=0\}$ with degree $d+2$ and $r:={\rm mdr}(f)$, namely
@@
where the left-hand side follows from the fact that for an ordinary singularity $p\in{\rm Sing}(C)$ of multiplicity $m_{p}<5$ one has $\tau_{p}=\mu_{p}=(m_{p}-1)^{2}$, see for instance [11, Section 2]. The above equation decodes the fact that if we can find a conic-line arrangement with ordinary double, ordinary triple, and ordinary quadruple intersection points with suitable $r$, then our arrangement is free. Based on this discussion, our last missing element is to find numerical constraints on $r={\rm mdr}(f)$. We have the following result.
@@
Let $\mathcal{CL}=\{f=0\}$ be an arrangement of $d\geqslant 3$ lines and one smooth conic that admits only ordinary singularities of multiplicity $<5$. Assume that $\mathcal{CL}$ is free, then
@@
Since $\mathcal{CL}$ is free, then by Theorem 2.4 one has ${\rm mdr}(f)\leqslant\lfloor\frac{d+1}{2}\rfloor$. Moreover, if $\mathcal{CL}$ admits only ordinary singularities with multiplicities $<5$, then $\alpha_{\mathcal{CL}}=\frac{1}{2}$, and this follows from Remark 2.7. This gives us that
@@
For $d\geqslant 3$ and
$r:=r(d)\in\bigg{\{}\bigg{\lceil}\frac{d-2}{2}\bigg{\rceil},\bigg{\lfloor}\frac{d+1}{2}\bigg{\rfloor}\bigg{\}},$ (4)
we consider the following Diophantine system of equations:
@@
Now our goal is to find all non-negative integer solutions $(n_{2},n_{3},n_{4})$ to (5), depending on $d$ and $r=r(d)$, to get a complete weak combinatorial description of the expected free conic-line arrangements. The last step boils down to deciding on the existence/non-existence of a geometric realization of a given weak combinatorics, which is a completely non-trivial problem. It is worth emphasizing here that if we can find an arrangement $\mathcal{CL}=\{f=0\}$ that satisfies conditions (5) with $r={\rm mdr}(f)$ that satisfies (4), then $\mathcal{CL}$ is automatically free. In our classification we use a well-known (even folkloric) result which tells us that all loopless matroids of rank $3$ with up to $6$ elements are vector matroids which can be represented geometrically as line arrangements in $\mathbb{P}^{2}_{\mathbb{K}}$ with $\mathbb{K}$ being any infinite field.
@@
    Since $\tau(\mathcal{CL}_{1})=12$, because $(n_{2},n_{3},n_{4})=(3,0,1)$, and ${\rm mdr}(Q_{1})=2$, hence both (4) and (5) are satisfied.
@@
    Since $(n_{2},n_{3},n_{4})=(0,3,0)$, $\tau(\mathcal{CL}_{2})=12$ and ${\rm mdr}(Q_{2})=2$, hence both (4) and (5) are satisfied.
@@
    Observe that $\tau(\mathcal{CL}_{3})=19$, since we have $(n_{2},n_{3},n_{4})=(2,2,1)$, and ${\rm mdr}(Q_{3})=2$, so both (4) and (5) are satisfied.
@@
    Since $\tau(\mathcal{CL}_{4})=28$, because we get the required intersections, and ${\rm mdr}(Q_{4})=2$, hence both (4) and (5) are satisfied.
@@
    Observe that $\tau(\mathcal{CL}_{5})=27$ and ${\rm mdr}(Q_{5})=3$, hence both (4) and (5) are satisfied
@@
    Observe that $\tau(\mathcal{CL}_{6})=39$, since $(n_{2},n_{3},n_{4})=(3,0,4)$, and ${\rm mdr}(Q_{6})=2$, hence both (4) and (5) are satisfied.
@@
    Observe that $\tau(\mathcal{CL}_{7})=37$, since $(n_{2},n_{3},n_{4})=(3,4,2)$, and ${\rm mdr}(Q_{7})=3$, hence both (4) and (5) are satisfied.
@@
    Observe that $\tau(\mathcal{CL}_{8})=37$, since $(n_{2},n_{3},n_{4})=(6,1,3)$, and ${\rm mdr}(Q_{8})=3$, hence both (4) and (5) are satisfied.
    To complete our classification for $d=6$, we need to show that the weak combinatorics $(n_{2},n_{3},n_{4})\in\{(0,3,3),(0,7,1)\}$ cannot be realized geometrically over the reals. In order to verify this claim we can start by looking at real line arrangements with only double, triple and quadruple points, and the following weak combinatorics are admissible:
    $(n_{2}^{\mathcal{L}},n_{3}^{\mathcal{L}},n_{4}^{\mathcal{L}})\in\{(15,0,0),(12,1,0),(9,2,0),(3,4,0),(9,0,1),(6,1,1)\},$
    where $n_{i}^{\mathcal{L}}$ denotes the number of $i$-fold points of a given line arrangement $\mathcal{L}\subset\mathbb{P}^{2}_{\mathbb{R}}$. If we add a smooth conic to a given line arrangement, then by Bézout’s Theorem the number of singularities of the resulting conic-line arrangement can either remain the same or increase. This implies that we can restrict our attention to the following weak combinatorics:
    $(n_{2}^{\mathcal{L}},n_{3}^{\mathcal{L}},n_{4}^{\mathcal{L}})\in\{(3,4,0),(6,1,1)\}.$
    Furthermore, it follows from the above argument that we cannot construct a conic-line arrangement with the weak-combinatorics $(n_{2},n_{3},n_{4})=(0,3,3)$ and we are left with the case $(n_{2},n_{3},n_{4})=(0,7,1)$. If we start with $\mathcal{L}$ such that $(n_{2}^{\mathcal{L}},n_{3}^{\mathcal{L}},n_{4}^{\mathcal{L}})=(3,4,0)$, then we have to add a conic that passes through one triple point and all double intersection points, but in that way we produce a conic-line arrangement with the weak combinatorics $(n_{2},n_{3},n_{4})=(3,6,1)$. Now we look at $(n_{2}^{\mathcal{L}},n_{3}^{\mathcal{L}},n_{4}^{\mathcal{L}})=(6,1,1)$. This arrangement is constructed in the following way. We take $3$ lines passing through a point $P\in\mathbb{P}^{2}_{\mathbb{R}}$, and $2$ lines passing through a point $Q\in\mathbb{P}^{2}_{\mathbb{R}}$ such that $P\neq Q$. Then we join these two pencils by the line passing through $P$ and $Q$. To construct our conic-line arrangement, we have to add a conic that would pass through $6$ double points of $\mathcal{L}$, but such a conic cannot exist by Bézout’s Theorem since in the line arrangement we can find $3$ collinear double points.
@@
    We can easily observe that $\tau(\mathcal{CL}_{9})=49$ and ${\rm mdr}(Q_{9})=3$, hence both (4) and (5) are satisfied.
@@
    One can easily check that $\tau(\mathcal{CL}_{10})=48$ and ${\rm mdr}(Q_{10})=4$, hence both (4) and (5) are satisfied.
@@
    One can check that $\tau(\mathcal{CL}_{11})=48$ and ${\rm mdr}(Q_{10})=4$, so both (4) and (5) are satisfied.
@@
    We can check that $(n_{2},n_{3},n_{4})=(5,5,4)$, so we have $\tau(\mathcal{CL}_{12})=61$, and ${\rm mdr}(Q_{12})=4$, so both (4) and (5) are satisfied.
@@
    We can check that $(n_{2},n_{3},n_{4})=(2,8,3)$, so we have $\tau(\mathcal{CL}_{13})=61$, and ${\rm mdr}(Q_{13})=4$, so both (4) and (5) are satisfied.
@@
    We can check that $(n_{2},n_{3},n_{4})=(8,2,5)$, so we have $\tau(\mathcal{CL}_{14})=61$, and ${\rm mdr}(Q_{14})=4$, so both (4) and (5) are satisfied.
@@
    Since $\tau(\mathcal{CL}_{15})=76$ and ${\rm mdr}(\mathcal{CL}_{15})=4$, hence both (4) and (5) are satisfied.
@@
    Since $\tau(\mathcal{CL}_{16})=93$ and ${\rm mdr}(Q_{16})=4$, so both (4) and (5) are satisfied.
@@
Remark 3.3.
As we can see, in the cases $d=7,8,9,10$ some weak combinatorics remained untouched by us since we could not find their geometric realization over the real numbers. It is a separate and quite difficult problem to decide whether these weak combinatorics can be realized over the real or the complex numbers.